OpenStudy (anonymous):

Simplify. 8 + 3 [3 - (1)^6] Please explain and can some please tell Me what [ < these mean

OpenStudy (dan815):

those are just brackets like () but just []

OpenStudy (anonymous):

But how do you simplify it if you know?

OpenStudy (anonymous):

So it doesn't matter if it's [] or ()???

OpenStudy (dan815):

okay just work with it like normal brackets do you know PEMDAS OR BEDMAS

OpenStudy (dan815):

ya it doesnt matter

OpenStudy (anonymous):

Yeah pemdas

OpenStudy (dan815):

8 + 3 (3 - (1)^6)

OpenStudy (dan815):

its just to help you see which brackets go where easier

OpenStudy (anonymous):

I'll do pemdas for this then?

OpenStudy (dan815):

yes

OpenStudy (anonymous):

With pemdas when it motions exponents what do I do with the exponents?

OpenStudy (anonymous):

Do I multiply them with the whole number?

OpenStudy (dan815):

8 + 3 (3 - (1)^6) , when 1^6= 1*1*1*1*1*1=1 8+3(3-1) 8+3(2) 8+6 14

OpenStudy (anonymous):

Thanksssss